ydx+2(y^4-x)dy=0


1
Expert's answer
2022-02-01T07:35:03-0500

The given equation can be written as "y~dx = -2(y^{4} - x)dy". It would be easier to solve this equation in terms of "\frac{dx}{dy}."


"\dfrac{dx}{dy} = -\dfrac{(2y^{4} - x)}{y} = -2y^3 + \dfrac{2x}{y}\\
\dfrac{dx}{dy} - \dfrac{2x}{y} = -2y^3 \\"

which is of the form

"\dfrac{dx}{dy} + P(y)x = Q(y)."

whose solution is given by "xe^{\int Pdy} = \int Q e^{\int Pdy} dy + c", where "e^{\int Pdy}"  is called the integrating factor.


Here "P(y) = -\frac{2}{y}" and "e^{\int Pdy} = e^{\int \frac{-2}{y}} = \frac{1}{y^{2}}". Hence the solution of the given equation is


"\dfrac{x}{y^{2}} = \int -2y^{3} \cdot \frac{1}{y^{2}}dy + c\\
\dfrac{x}{y^{2}} = \int -2y dy + c\\
\dfrac{x}{y^{2}} = -y^{2} + c\\
\therefore x = cy^{2} - y^{4}"